VEGAS Team Meeting

434-817-6443# (telephone)

Attendees: Ray, Adam, Joe, Glenn, Sam, Mark, Paul, Marty, Jason, Dan, John, Dave


  1. Baseline tests, results and actions
  2. 'Spike' tests, results and actions
  3. Next meeting?

Meeting notes:

  1. Baseline tests, results and actions
    • After digital filtering the gain difference was such that re-quantized values were in error
    • Adding a gain adjustment cleared problem
    • Consider adding it as a dynamic balance adjustment in the manager based on an algorithm that Anish has created (Modes 10-19)
      • Anish to draft MR
    • With the new levels the baseline issues are much better
    • DONE Agreed
  2. 'Spike' tests, results and actions
    • Details and test data can be seen here.
    • Next steps
      • Try the shift setting tests recommended (ffff v. aaaa)
        • shifts every stage, which creates quantization effects due to the small numbers
        • may end up with different shift settings for bands depending on amount of FRI
    • Recheck OGP(?) calibrations on the next maintenance day
    • First fix build environment (NRAO & UCB)
    • Then look for Mode 2&3 spikes/fix spikes
    • Then look at spikes in Mode 4 (believe problem is different)
  3. Next meeting
    • Tuesday afternoon for updates

-- MartyBloss - 2014-08-29
Topic revision: r2 - 2014-08-29, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback